Bobby is 3years older than Chris last year the sum of their age was 39,how old each now
Found 2 solutions by stanbon, TeachMath:
Answer by stanbon(75887)

Bobby is 3years older than Chris last year the sum of their age was 39,how old each now
-----
Equations:
B = C+3
B+C = 39
-----
Substitute for "B" and solve for "C"::
C+3 + C = 39
2C = 36
Chris = 18 yrs.
----
Solve for "B"::
Bobby = C+3 = 18+3 = 21 yrs
-------------
Cheers,
Stan H.

Answer by TeachMath(96)

The dude is WRONG!!

Chris is 19, and Bobby is 22